Grade 12th passMechanics

  1. ...........Prove that v =w(omaga) × r
  2. ............quickanswer

Biplob Mondol
7 Years agoGrade 12th pass

1 Answer

Sai Sanjana Gedda
7 Years ago
As we know theta=arc length/radius =L/r                                                                                                   L=r*theta                                                                                                                                                velocity=displacement/time taken                                                                                                     v=dl/dt (for a small movement of the particle)
from above …...........                                                                                                                             v=d(r*theta)/dt                                                                                                                                       as r is constant it comes out                                                                                                                  =r(d*theta/dt)                                                                                                                                         note: w(omega)=d*theta/dt        …........ here w is angular velocity=angular displacement/time taken ….....so v=rw